OpenStudy (anonymous):

write expression as a sum or difference of logarithms. express power as factors. log base 5 ((x^6)/(x-6)) x >6

jimthompson5910 (jim_thompson5910):

you would use the rule log(A/B) = log(A) - log(B)

jimthompson5910 (jim_thompson5910):

and this rule as well log(A^B) = B*log(A)

OpenStudy (anonymous):

okay so it will look like log (base5)(x^6)-log (base5)(x-6) (x)>6

jimthompson5910 (jim_thompson5910):

ignore the x > 6 part

jimthompson5910 (jim_thompson5910):

that doesn't need to be written again

jimthompson5910 (jim_thompson5910):

it's just there to guarantee that the expression (x^6)/(x-6) is always positive

jimthompson5910 (jim_thompson5910):

after getting that, use the rule log(A^B) = B*log(A)

OpenStudy (anonymous):

so it will look like (6log(x))-log(x-6)

jimthompson5910 (jim_thompson5910):

good, each log is assumed to be base 5, but no need to write "base 5" for each log if you make that statement

OpenStudy (anonymous):

okay so that will be the simplified form

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw